Usin the method of false position. Find the root of equation x^6 - x^4 - x^3 - 1 = 0
1
Expert's answer
2020-10-25T19:15:35-0400
"x^6-x^4-x^3-1=x^4(x^2-1)-(x^3+1)="

"=x^4(x-1)(x+1)-(x+1)(x^2-x+1)="

"=(x-1)(x^5+x^4-x^2+x-1)"

"x1=1"


Given "f(x)=x^6-x^4-x^3-1"

Find points "a" and "b"  such that "a<b"  and  "f(a)f(b)<0"

Let "a=1, b=1.5"

"f(a)=f(1)=1^6-1^4-1^3-1=-2"

"f(b)=f(1.5)=1.5^6-1.5^4-1.5^3-1=1.953125"

"f(a)f(b)=f(1)f(1.5)<0"

Using iterative Regula-Falsi Formula, the first approximation,


"x_1=a-\dfrac{b-a}{f(b)-f(a)}f(a)=""=1-\dfrac{1.5-1}{1.953125+2}(-2)\approx1.252964"



"f(1.252964)=-1.562404"



"x_2=1.252964-\dfrac{1.5-1.252964}{1.953125+1.562404}(-1.562404)\approx""\approx1.362754"


"f(1.362754)=-0.574795"



"x_3=1.362754-\dfrac{1.5-1.362754}{1.953125+0.574795}(-0.574795)\approx""\approx1.393961"


"f(1.393961)=-0.147639"



"x_4=1.393961-\dfrac{1.5-1.393961}{1.953125+0.147639}(-0.147639)\approx""\approx1.401413"


"f(1.401413)=-0.034199"



"x_5=1.401413-\dfrac{1.5-1.401413}{1.953125+0.034199}(-0.034199)\approx""\approx1.403110"


"f(1.403110)=-0.007729"



"x_6=1.403110-\dfrac{1.5-1.403110}{1.953125+0.007729}(-0.007729)\approx""\approx1.403492"



"f(1.403492)=-0.000390"


"x_7=1.403492-\dfrac{1.5-1.403492}{1.953125+0.000390}(-0.000390)\approx""\approx1.403597"


"f(1.403597)=-0.000087"


"x_8=1.403597-\dfrac{1.5-1.403597}{1.953125+0.000087}(-0.000087)\approx""\approx1.403601"


"f(1.403601)=-0.000020"



"x_8=1.403601-\dfrac{1.5-1.403601}{1.953125+0.000020}(-0.000020)\approx""\approx1.403602"



"f(1.403602)=-0.000002"


If "\varepsilon=0.0001," then "|f(x)|<0.0001"

Root of "f(x)" is "x=1.403602"